OpenStudy (anonymous):

prove siny +tany/over/ 1+secy=siny

OpenStudy (psymon):

\[\frac{ siny+tany }{ 1+secy }=siny\]?

OpenStudy (anonymous):

yes

OpenStudy (psymon):

\[\frac{ siny+\frac{ siny }{ cosy } }{ 1+\frac{ 1 }{ cosy } }\] Make those into common denominators and simplify it out. It'll gert your answer : )

OpenStudy (anonymous):

oh but i seem to be stuck on this part sinycosy=siny/cosy large fraction 1=cosy/cosy

OpenStudy (psymon):

Alrighty, lemme see : )

OpenStudy (anonymous):

I made a mistake the equal signs should be + sign

OpenStudy (psymon):

\[\frac{ siny + \frac{ siny }{ cosy } }{ 1+\frac{ 1 }{ cosy } }\]Common denominator makes this: \[\frac{ \frac{ sinycosy+siny }{ cosy } }{ \frac{ cosy+1 }{ cosy } }\]I can factor out the sin on top to have: \[\frac{ \frac{ \sin(cosy+1) }{ cosy } }{ \frac{ cosy+1 }{ cosy } }\] Can ya see where this will lead now? :P

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

I cant really tell how you are seeting them up

OpenStudy (anonymous):

do you think you could draw the equation

OpenStudy (psymon):

Yeah, sure.

OpenStudy (anonymous):

thank you

OpenStudy (psymon):

 Now combining into one fraction I get: